# Harmony in Divergence: Towards Fast, Accurate, and Memory-efficient Zeroth-order LLM Fine-tuning

### v1

@@ D.2 Llama Experiments
FT 94.2 81.2 91.4 72.2 63.8 65.8 78.2 79.6 40.3
MeZO 88.8 67.4 73.2 78.0 56.6 63.4 64.8 61.9 27.8
DiZO 90.0 68.2 76.7 76.8 57.8 63.8 64.2 63.2 29.7
Table 13: Experiments results on Llama3-8B for seven classification datasets and two text generation datasets (with 1000 training samples). Better results between MeZO and DiZO are highlighted in bold.
Task Task Type SST-2 RTE CB WSC SQuAD
@@
MeZO 90.0 67.8 71.4 60.2 67.0
DiZO 91.5 69.4 73.2 63.4 67.4
Appendix E Proof
We consider a neural network with $L$ layers (or parameter blocks) and wish to estimate the gradient of some loss function $\mathcal{L}(\boldsymbol{\theta};\mathcal{B})$ with respect to all parameters $\boldsymbol{\theta}$. We use a two-point finite-difference (zero-order) method with directions drawn from an isotropic distribution. We show below why the expected norm-squared of the resulting gradient estimator is identical (or follows the same dimension-based law) for each layer/block.
Consider the $\ell$-th layer. Its estimator is
$\widehat{\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}}\;=\;\frac{1}{q}\,\sum_{i=1}^{q}\Bigl{[}\,\underbrace{\frac{\mathcal{L}\!\bigl{(}\boldsymbol{\theta}+\epsilon\boldsymbol{u}_{i}\bigr{)}\;-\;\mathcal{L}\!\bigl{(}\boldsymbol{\theta}-\epsilon\boldsymbol{u}_{i}\bigr{)}}{2\,\epsilon}}_{\displaystyle\Delta_{i}}\Bigr{]}\,\boldsymbol{u}_{i}^{(\ell)},$
where $\Delta_{i}$ is the same scalar for all blocks. We want
$\mathbb{E}\Bigl{[}\|\widehat{\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}}\|^{2}\Bigr{]}.$
Note that:
@@ Appendix E Proof
    $\Delta_{i}$ does not depend on $\ell$; it is a single scalar for each direction $i$.
@@
    $\boldsymbol{u}_{i}^{(\ell)}$ is the sub-vector of $\boldsymbol{u}_{i}$ associated to the $\ell$-th block.
@@
    $\boldsymbol{u}_{i}$ is drawn from an isotropic distribution in $\mathbb{R}^{d}$, meaning each coordinate has zero mean, unit variance, and there is no cross-correlation between different coordinates. Thus, different blocks $\boldsymbol{u}_{i}^{(\ell)}$ and $\boldsymbol{u}_{i}^{(m)}$ (for $\ell\neq m$) are uncorrelated, and each block $\boldsymbol{u}_{i}^{(\ell)}$ has an identity covariance in its own subspace $\mathbb{R}^{d_{\ell}}$.
Hence, when we expand
@@
$\|\widehat{\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}}\|^{2}\;=\;\Bigl{\|}\frac{1}{q}\,\sum_{i=1}^{q}\Delta_{i}\,\boldsymbol{u}_{i}^{(\ell)}\Bigr{\|}^{2},$
the expectation w.r.t. $\{\boldsymbol{u}_{i}\}$ depends on $\ell$ only through the dimension $d_{\ell}$, not through any other distributional asymmetry. If $d_{\ell}$ are the same for all $\ell$, then the second moment is literally the same across all blocks. If $d_{\ell}$ differ, the dependence is only a (known) function of $d_{\ell}$.
In short, isotropy ensures that
$\mathbb{E}\!\bigl{[}\|\widehat{\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}}\|^{2}\bigr{]}\quad\text{is the same functional form of }\|\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}\|^{2}\text{ for each layer }\ell.$
Therefore, in the simplest scenario where $d_{\ell}$ are all the same, each layer gets the same second-moment behavior for its gradient estimator.
Appendix F Implementation

### v3

@@ E.4 Llama Fine-tuning
FT 94.2 (0.4) 81.2 (2.1) 91.4 (4.7) 72.2 (4.2) 63.8 (1.8) 65.8 (2.3) 78.2 (3.2) 79.6 (2.9) 40.3 (1.2)
MeZO 88.8 (1.1) 67.4 (1.7) 73.2 (2.4) 78.0 (4.4) 56.6 (3.8) 63.4 (2.3) 64.8 (3.1) 61.9 (2.7) 27.8 (2.0)
HiZOO 89.5 (1.4) 67.1 (1.3) 74.4 (1.9) 78.8 (4.7) 56.3 (3.1) 64.4 (2.4) 64.3 (2.9) 61.7 (2.8) 28.6 (3.0)
DiZO 90.0 (0.9) 68.2 (1.6) 76.7 (3.3) 76.8 (3.8) 57.8 (4.2) 63.8 (1.7) 64.2 (2.9) 63.2 (2.7) 29.7 (1.3)
Table E.6: Experiments results on Llama3-8B for seven classification datasets and two text generation datasets (with 1000 training samples). Better results between MeZO and DiZO are highlighted in bold.
Task
Task Type
SST-2 RTE CB WSC SQuAD
@@
MeZO 90.0 (0.7) 67.8 (1.4) 71.4 (2.2) 60.2 (1.6) 67.0 (2.6)
HiZOO 91.1 (0.9) 68.2 (1.3) 71.4 (2.9) 62.2 (1.9) 68.3 (3.1)
DiZO 91.5 (0.8) 69.4 (1.8) 73.2 (3.1) 63.4 (2.9) 67.4 (2.1)
E.5 Fine-tuning on MMLU and MT-Bench
To demonstrate the generalizability of DiZO in more realistic and challenging scenarios, we evaluate our method on MMLU and MT-Bench benchmarks. we follow the setting in [31, 32], fine-tune on the Alpaca GPT-4 dataset [33], which consists of 52k conversations, and then evaluate. We conduct experiments based on Llama2-7B and Llama3-8B, the results are shown in Table E.7 and Table E.8, respectively.
Table E.7: Results of fine-tuning Llama2-7B on more challenging benchmarks, better results are highlighted in bold.
MT-Bench MMLU (5 shot) GPU hours
Zero-shot 3.93 45.87 -
MeZO 4.59 45.22 100%
HiZOO 4.62 45.42 92%
DiZO 4.79 45.91 78%
Table E.8: Results of fine-tuning Llama3-8B on more challenging benchmarks, better results are highlighted in bold.
MT-Bench MMLU (5 shot) GPU hours
Zero-shot 5.46 65.20 -
MeZO 5.89 65.08 100%
HiZOO 5.93 65.20 95%
DiZO 6.15 65.42 83%
E.6 Compare with Sparse Technique
We conducted a direct comparison between our DiZO and Sparse MeZO [13] in terms of both accuracy and training efficiency across two datasets and two model sizes. The results are presented in Table R.6 and Table R.7. From the accuracy perspective, DiZO consistently outperforms Sparse MeZO under all evaluated settings, demonstrating the effectiveness of our projection-based approach.
From the efficiency perspective, Sparse MeZO requires generating the sparsity mask dynamically during training. Compared to DiZO, Sparse MeZO requires longer GPU hours for about 20%, and slows the throughput for more than 30%. Moreover, with the grows of model size, the throughput of Sparse MeZO will further decrease, due to the growing cost of maintaining and updating the sparse mask. While Sparse MeZO reduces the number of training iterations, its lower throughput results in longer total GPU hours compared to DiZO. These findings demonstrate that DiZO not only delivers better accuracy but also achieves more practical training efficiency compared to Sparse MeZO.
Table E.9: Acc and Speed Comparison on OPT-2.7B.
Method Dataset Accuracy Throughput #Train Iter. GPU Hours
MeZO SST2 90.0 3.3it/s 100% 100%
Sparse MeZO SST2 91.4 2.3it/s 55% 79%
DiZO SST2 92.3 1.9it/s 52% 56%
MeZO RTE 63.5 1.7it/s 100% 100%
Sparse MeZO RTE 67.1 1.1it/s 50% 73%
DiZO RTE 68.4 1.5it/s 60% 62%
Table E.10: Acc and Speed Comparison on OPT-6.7B.
Method Dataset Accuracy Throughput #Train Iter. GPU Hours
MeZO SST2 90.2 1.8it/s 100% 100%
Sparse MeZO SST2 91.9 1.0it/s 47% 84%
DiZO SST2 92.4 1.7it/s 62% 69%
MeZO RTE 73.2 0.6it/s 100% 100%
Sparse MeZO RTE 74.3 0.3it/s 39% 88%
DiZO RTE 74.8 0.5it/s 65% 81%
Appendix F Theoretical Analysis
F.1 Variance Symmetry under Isotropic Perturbations
We consider a neural network with $L$ layers (or parameter blocks) and analyze the zeroth-order gradient estimator constructed via two-point finite differences. Let $\mathcal{L}(\boldsymbol{\theta};\mathcal{B})$ denote the loss evaluated on mini-batch $\mathcal{B}$, with $\boldsymbol{\theta}=(\boldsymbol{\theta}^{(1)},\dots,\boldsymbol{\theta}^{(L)})$ the full parameter vector. To estimate the gradient $\nabla\mathcal{L}$ without access to derivatives, we apply a two-sided estimator along randomly sampled directions $\boldsymbol{u}_{i}\in\mathbb{R}^{d}$:
$\widehat{\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}}\;=\;\frac{1}{q}\,\sum_{i=1}^{q}\underbrace{\frac{\mathcal{L}(\boldsymbol{\theta}+\epsilon\boldsymbol{u}_{i})-\mathcal{L}(\boldsymbol{\theta}-\epsilon\boldsymbol{u}_{i})}{2\epsilon}}_{\Delta_{i}}\,\boldsymbol{u}_{i}^{(\ell)},$
where $\boldsymbol{u}_{i}^{(\ell)}$ is the sub-vector of direction $\boldsymbol{u}_{i}$ corresponding to layer $\ell$.
We aim to characterize the variance of this estimator, in particular:
$\mathbb{E}\left[\left\|\widehat{\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}}\right\|^{2}\right].$
Key assumptions:
@@ F.1 Variance Symmetry under Isotropic Perturbations
    Each $\boldsymbol{u}_{i}$ is drawn independently from an isotropic distribution in $\mathbb{R}^{d}$, i.e., $\mathbb{E}[\boldsymbol{u}_{i}\boldsymbol{u}_{i}^{\top}]=I$.
@@
    The scalar $\Delta_{i}$ is the same across all parameter blocks for a given $i$, as it depends only on the global perturbation.
@@
    Each block $\boldsymbol{u}_{i}^{(\ell)}$ has zero mean, unit covariance in its own subspace $\mathbb{R}^{d_{\ell}}$, and is uncorrelated with other blocks $\boldsymbol{u}_{i}^{(m)}$ for $\ell\neq m$.
Variance Expansion:
We examine the norm-squared of the estimator:
@@
$\left\|\widehat{\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}}\right\|^{2}=\left\|\frac{1}{q}\sum_{i=1}^{q}\Delta_{i}\boldsymbol{u}_{i}^{(\ell)}\right\|^{2}.$
Taking expectation over $\{\boldsymbol{u}_{i}\}$:
$\mathbb{E}\left[\left\|\widehat{\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}}\right\|^{2}\right]=\frac{1}{q^{2}}\sum_{i=1}^{q}\mathbb{E}\left[\Delta_{i}^{2}\|\boldsymbol{u}_{i}^{(\ell)}\|^{2}\right]+\frac{1}{q^{2}}\sum_{i\neq j}\mathbb{E}\left[\Delta_{i}\Delta_{j}\langle\boldsymbol{u}_{i}^{(\ell)},\boldsymbol{u}_{j}^{(\ell)}\rangle\right].$
Due to independence and zero-mean isotropy, cross terms vanish, and we obtain:
$\mathbb{E}\left[\left\|\widehat{\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}}\right\|^{2}\right]=\frac{1}{q}\mathbb{E}\left[\Delta^{2}\cdot\|\boldsymbol{u}^{(\ell)}\|^{2}\right],$
where $\Delta$ and $\boldsymbol{u}^{(\ell)}$ are representative samples from the same distribution.
Conclusion.
Since $\Delta$ is shared across layers and $\boldsymbol{u}^{(\ell)}$ has expected squared norm proportional to the layer dimension $d_{\ell}$, we conclude:
$\mathbb{E}\left[\left\|\widehat{\nabla_{\boldsymbol{\theta}^{(\ell)}}\mathcal{L}}\right\|^{2}\right]\propto d_{\ell}.$
In other words, the second-moment of the gradient estimator depends on the layer only through its dimensionality $d_{\ell}$, and not through any asymmetry in the distribution of direction vectors. If $d_{\ell}$ are the same for all $\ell$, each block exhibits identical expected variance.
This property justifies using uniform per-layer treatment in analysis and initialization when random direction sampling is isotropic.
F.2 The Proof of Convergence Analysis
We now prove Theorem 1.
Proof.
By $L_{f}$-smoothness (Assumption 5.1), for any $\bm{\theta},\bm{\theta}^{\prime}$:
$\mathcal{L}(\bm{\theta}^{\prime})\;\leq\;\mathcal{L}(\bm{\theta})\;+\;\langle\nabla\mathcal{L}(\bm{\theta}),\,\bm{\theta}^{\prime}-\bm{\theta}\rangle\;+\;\frac{L_{f}}{2}\,\|\bm{\theta}^{\prime}-\bm{\theta}\|^{2}.$
Set $\bm{\theta}=\bm{\theta}_{t}$ and $\bm{\theta}^{\prime}=\bm{\theta}_{t+1}$. We get:
$\mathcal{L}(\bm{\theta}_{t+1})\;\leq\;\mathcal{L}(\bm{\theta}_{t})\;+\;\bigl\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,\bm{\theta}_{t+1}-\bm{\theta}_{t}\bigr\rangle\;+\;\frac{L_{f}}{2}\,\|\bm{\theta}_{t+1}-\bm{\theta}_{t}\|^{2}.$
Taking conditional expectation $\mathbb{E}_{t}[\cdot]\!:=\!\mathbb{E}[\cdot\mid\bm{\theta}_{t}]$ yields
$\mathbb{E}_{t}\bigl[\mathcal{L}(\bm{\theta}_{t+1})-\mathcal{L}(\bm{\theta}_{t})\bigr]\;\leq\;\mathbb{E}_{t}\bigl[\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,\bm{\theta}_{t+1}-\bm{\theta}_{t}\rangle\bigr]\;+\;\frac{L_{f}}{2}\,\mathbb{E}_{t}\bigl[\|\bm{\theta}_{t+1}-\bm{\theta}_{t}\|^{2}\bigr].$ (8)
Denote $\widetilde{\bm{\theta}}_{t+1}=\bm{\theta}_{t}-\eta\,g_{t}$. Then
$\bm{\theta}_{t+1}\;=\;\text{Proj}_{\mathcal{S}}(\widetilde{\bm{\theta}}_{t+1}),\qquad\bm{\theta}_{t+1}-\bm{\theta}_{t}\;=\;(\bm{\theta}_{t+1}-\widetilde{\bm{\theta}}_{t+1})\;-\;\eta\,g_{t}.$
Hence
$\bigl\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,\bm{\theta}_{t+1}-\bm{\theta}_{t}\bigr\rangle\;=\;\bigl\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,\bm{\theta}_{t+1}-\widetilde{\bm{\theta}}_{t+1}\bigr\rangle\;-\;\eta\,\bigl\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,g_{t}\bigr\rangle.$
Because $\bm{\theta}_{t+1}$ is the nearest point in $\mathcal{S}$ to $\widetilde{\bm{\theta}}_{t+1}$ (by definition of projection), we have
$\|\bm{\theta}_{t+1}-\widetilde{\bm{\theta}}_{t+1}\|\;\leq\;\|\bm{\theta}_{t}-\widetilde{\bm{\theta}}_{t+1}\|\;=\;\eta\,\|g_{t}\|.$
Thus
$\|\bm{\theta}_{t+1}-\widetilde{\bm{\theta}}_{t+1}\|\;\leq\;\eta\,\|g_{t}\|.$ (9)
Taking conditional expectation:
$\mathbb{E}_{t}\bigl[\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,\bm{\theta}_{t+1}-\bm{\theta}_{t}\rangle\bigr]\;=\;\mathbb{E}_{t}\bigl[\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,\bm{\theta}_{t+1}-\widetilde{\bm{\theta}}_{t+1}\rangle\bigr]\;-\;\eta\,\mathbb{E}_{t}\bigl[\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,g_{t}\rangle\bigr].$
Using (9) with Cauchy–Schwarz:
$\bigl|\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,\bm{\theta}_{t+1}-\widetilde{\bm{\theta}}_{t+1}\rangle\bigr|\;\leq\;\|\nabla\mathcal{L}(\bm{\theta}_{t})\|\;\|\bm{\theta}_{t+1}-\widetilde{\bm{\theta}}_{t+1}\|\;\leq\;\eta\,\|\nabla\mathcal{L}(\bm{\theta}_{t})\|\;\|g_{t}\|.$
Moreover, the two-point estimator is unbiased, so
$\mathbb{E}_{t}\bigl[\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,g_{t}\rangle\bigr]\;=\;\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,\mathbb{E}_{t}[g_{t}]\rangle\;=\;\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}.$
Hence
$\mathbb{E}_{t}\bigl[\langle\nabla\mathcal{L}(\bm{\theta}_{t}),\,\bm{\theta}_{t+1}-\bm{\theta}_{t}\rangle\bigr]\;\leq\;\eta\,\|\nabla\mathcal{L}(\bm{\theta}_{t})\|\;\mathbb{E}_{t}\bigl[\|g_{t}\|\bigr]\;-\;\eta\,\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}.$ (10)
Again from the projection property:
$\|\bm{\theta}_{t+1}-\bm{\theta}_{t}\|\;=\;\|\text{Proj}_{\mathcal{S}}(\widetilde{\bm{\theta}}_{t+1})-\bm{\theta}_{t}\|\;\leq\;\|\widetilde{\bm{\theta}}_{t+1}-\bm{\theta}_{t}\|\;=\;\eta\,\|g_{t}\|.$
Thus
$\|\bm{\theta}_{t+1}-\bm{\theta}_{t}\|^{2}\;\leq\;\eta^{2}\,\|g_{t}\|^{2}.$
Taking expectation completes:
$\mathbb{E}_{t}\bigl[\|\bm{\theta}_{t+1}-\bm{\theta}_{t}\|^{2}\bigr]\;\leq\;\eta^{2}\,\mathbb{E}_{t}\bigl[\|g_{t}\|^{2}\bigr].$
Substitute (10) and the above bound into (8):
$\mathbb{E}_{t}\bigl[\mathcal{L}(\bm{\theta}_{t+1})-\mathcal{L}(\bm{\theta}_{t})\bigr]\;\leq\;\eta\,\|\nabla\mathcal{L}(\bm{\theta}_{t})\|\;\mathbb{E}_{t}[\|g_{t}\|]\;-\;\eta\,\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}\;+\;\frac{L_{f}}{2}\,\eta^{2}\,\mathbb{E}_{t}[\|g_{t}\|^{2}].$
Taking total expectation and summing over $t=0$ to T1T-1,
$\mathbb{E}\bigl[\mathcal{L}(\bm{\theta}_{T})\bigr]\;-\;\mathcal{L}(\bm{\theta}_{0})\;\leq\;\sum_{t=0}^{T-1}\Bigl\{\eta\,\mathbb{E}\bigl[\|\nabla\mathcal{L}(\bm{\theta}_{t})\|\;\|g_{t}\|\bigr]\;-\;\eta\,\mathbb{E}\bigl[\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}\bigr]\;+\;\tfrac{L_{f}}{2}\,\eta^{2}\,\mathbb{E}\bigl[\|g_{t}\|^{2}\bigr]\Bigr\}.$
Assume that the two-point finite-difference estimator satisfies
$\mathbb{E}\bigl[\|g_{t}-\nabla\mathcal{L}(\bm{\bm{\theta}}_{t})\|^{2}\bigr]\;\leq\;\sigma^{2}\bigl(D,q\bigr),$
where $\sigma^{2}(D,q)$ grows primarily with $D$ (the maximum layer dimension) and the number of queries $q$, rather than the full sum of dimensions.
Therefore, we have
$\mathbb{E}\bigl[\|g_{t}\|^{2}\bigr]\;\leq\;c_{1}\,\bigl(\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\sigma^{2}(D,q)\bigr)$
for some constant $c_{1}>0$. In addition, $\mathbb{E}[\|\nabla\mathcal{L}(\bm{\theta}_{t})\|\;\|g_{t}\|]\leq\sqrt{\,\mathbb{E}[\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}]\;\mathbb{E}[\|g_{t}\|^{2}]}\,\leq\,c_{2}\,(\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}+\sigma^{2}(D,q))$ (for a constant $c_{2}$). Collecting terms and choosing $\eta=c/\sqrt{T}$ with $c>0$ sufficiently small ensures that the $\eta\,\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}$ term dominates the positive terms for large $T$. A standard telescoping argument shows
$\frac{1}{T}\,\sum_{t=0}^{T-1}\mathbb{E}\bigl[\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}\bigr]\;=\;O\!\Bigl(\frac{\sqrt{D}}{\sqrt{T}}\Bigr),$
which implies the stationarity measure
$\min_{0\leq t<T}\mathbb{E}\Bigl[\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}\Bigr]\;=\;O\!\Bigl(\frac{\sqrt{D}}{\sqrt{T}}\Bigr).$
This completes the proof. ∎
F.3 $\tau$-Stability of the Clipping Step
Assumptions.
  • A1.
    $L:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is $L_{f}$-smooth.
  • A2.
    Zeroth-order updates use step-size $\eta_{t}=\eta$ and produce an intermediate point $\theta_{t}^{\ell}=\theta_{t}-\eta g_{t}$ with an unbiased estimator $g_{t}$.
  • A3.
    After every $k$ iterations we apply projection clipping:
    $\theta_{t+1}^{(\ell)}=\theta_{t}^{(0)}+\rho_{t}^{(\ell)}\Delta\theta_{t}^{(\ell)},\quad\rho_{t}^{(\ell)}:=\text{clip}\big[\rho_{t}^{(\ell)},\,1-\tau,\,1+\tau\big],$
    where $\rho_{t}^{(\ell)}=\gamma_{t}^{(\ell)}/\|\Delta\theta_{t}^{(\ell)}\|_{2}$, and $\tau\in(0,1)$ is the scalar clipping width.
Let
$R_{\max}:=\max_{t,\ell}\|\Delta\theta_{t}^{(\ell)}\|_{2},\qquad G_{\max}:=\max_{t}\|g_{t}\|_{2}.$
Then for any $\tau$, we have the stability bound:
$\|\theta_{t+1}-\theta_{t}^{\ell}\|_{2}\leq\eta G_{\max}+\tau R_{\max}.$ (11)
Moreover, if $\tau$ satisfies
$\tau\leq c_{\tau}\frac{\eta G_{\max}}{R_{\max}},\quad 0<c_{\tau}\leq 1,$ (12)
then under the same conditions as Theorem 5.3, the projected DiZO iterates satisfy
$\min_{0\leq t<T}\mathbb{E}\big[\|\nabla L(\theta_{t})\|_{2}^{2}\big]=\mathcal{O}\!\left(\frac{\sqrt{D}}{\sqrt{T}}\right).$ (13)
That is, the original non-convex convergence rate is preserved. If $\tau$ violates (12), the bound inflates linearly with $\tau$, i.e.,
$\widetilde{\mathcal{O}}\!\big(\sqrt{D/T}+\tau/\eta\big).$
Proof.
For any $L_{f}$-smooth loss, we have
$L(\theta_{t+1})\leq L(\theta_{t})+\langle\nabla L(\theta_{t}),\theta_{t+1}-\theta_{t}\rangle+\frac{L_{f}}{2}\|\theta_{t+1}-\theta_{t}\|_{2}^{2}.$ (14)
Because $\theta_{t+1}$ is the nearest feasible point to $\theta_{t}^{\ell}$ (Euclidean projection), the proof in the paper shows
$\|\theta_{t}-\theta_{t}^{\ell}\|_{2}\leq\eta\|g_{t}\|_{2}.$ (15)
The additional clipping in $\rho_{t}^{(\ell)}$ imposes another bound:
$\|\theta_{t+1}^{(\ell)}-\theta_{t}^{(\ell)}\|_{2}=|\rho_{t}^{(\ell)}-1|\|\Delta\theta_{t}^{(\ell)}\|_{2}\leq\tau R_{\max},$ (16)
hence combining gives
$\|\theta_{t+1}-\theta_{t}\|_{2}\leq\eta G_{\max}+\tau R_{\max}.$ (17)
Plugging (17) into (14) and taking conditional expectation yields
$\mathbb{E}_{t}[L(\theta_{t+1})-L(\theta_{t})]\leq-\eta\mathbb{E}_{t}[\langle\nabla L(\theta_{t}),g_{t}\rangle]+L_{f}(\eta G_{\max}+\tau R_{\max})\eta G_{\max}.$
Using standard ZO analysis, the inner product term satisfies
$\frac{\eta}{2}\|\nabla L(\theta_{t})\|_{2}^{2}+\frac{\eta}{2}\sigma^{2},$
and after telescoping the left-hand side for $t=0,\dots,T-1$, dividing by $\eta T$, we obtain
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\!\left[\|\nabla L(\theta_{t})\|_{2}^{2}\right]\leq\mathcal{O}\!\left(\frac{\sqrt{D}}{\sqrt{T}}\right)+\mathcal{O}\!\left(\frac{\tau R_{\max}}{\eta}\right).$
If $\tau$ satisfies condition (12), the second term is dominated, yielding (13).
Remark.
Equation (12) implies that the projection impulse matches a single ZO step:
$\tau\approx\frac{\eta\|g_{t}\|_{2}}{\|\Delta\theta_{t}\|_{2}}.$
During fine-tuning, we typically observe $\|g_{t}\|_{2}/\|\Delta\theta_{t}\|_{2}\in[0.5,1.5]$ after warm-up, and learning rates $\eta\!\sim\!10^{-2}\!\text{–}\!10^{-1}$. This leads to $\tau$ in the range 0.050.050.300.30, with $\tau\!\approx\!0.2$ being the empirically stable choice across tasks.
Appendix G Implementation